OpenStudy (anonymous):

Eliminate the parameter t from the parametric equations. x=-1+sint, y=3+cost

ganeshie8 (ganeshie8):

hint : \(\sin^2t + \cos^2t = 1\)

OpenStudy (anonymous):

i still cant get t eliminated

OpenStudy (anonymous):

solve for sin t and for cos t then square and add

OpenStudy (anonymous):

getting a high number which isn't possible

OpenStudy (anonymous):

\[(x+1)^2 + (y-3)^2 = 1\]

OpenStudy (anonymous):

circle with radius 1 centered at (-1, 3)

OpenStudy (anonymous):

i keep getting 6-dont know

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

\[x=-1+\sin t \Rightarrow \sin t = x + 1\]\[y=3+\cos t \Rightarrow \cos t = y -3\]

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

do i have to solve? this is something we havent learned yet, just ec

OpenStudy (anonymous):

you have to do what you have to do...

OpenStudy (anonymous):

k

OpenStudy (anonymous):

remember \[\sin^2 t + \cos^2 t = 1\]then substitue in and no more t!

OpenStudy (anonymous):

(2x+2) + (2y-9)=1...y=5 x=-1

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i guess i am confused

OpenStudy (anonymous):

after eliminating t, i thought i needed to solve further?

OpenStudy (anonymous):

\[\sin t = (x+1) \Rightarrow \sin^2 t = (x+1)^2\]

OpenStudy (anonymous):

likewise for cos t and then you have the curve in terms of x and y only, not t

OpenStudy (anonymous):

ok, so once t is removed there is no need to solve further?

OpenStudy (anonymous):

yeah, you look to see what you have... if it's a curve you recognize (hopefully).

OpenStudy (anonymous):

ah, ok

OpenStudy (anonymous):

\[(x+1)^2 + (y-3)^2 = 1\] is a circle with radius 1 , centered at (-1, 3)